OpenStudy (anonymous):

Which expression has a value of 15, using only four 4s and any of the following: + , – , · , ÷, exponents, and/or parentheses?

OpenStudy (mathstudent55):

Are there choices, or do you need to come with an expression?

OpenStudy (anonymous):

A. 44 – (4 ÷ 4) B. (4 · 4) – (4 + 4) C. (4 · 4) – (4 ÷ 4) D. (4 + 4 + 4) ÷ 4

OpenStudy (mathstudent55):

Evaluate each choice. Remember to do parentheses first. A. 44 – (4 ÷ 4) = 44 - 1 = 43 This is not it. B. (4 · 4) – (4 + 4) = 16 - 8 = 8 This is also not it. Can you evaluate C and D?

OpenStudy (anonymous):

i dont know im horrible at math